h(x)=-2x+11 Evaluate the function when x=-3,0 and 2.
Answer by oscargut(2103)

h(x)=-2x+11 Evaluate the function when x=-3,0 and 2
h(-3)=-2(-3)+11=17
h(2)=-2(2)+11=7